6. Show that for any sequence of events (F)y-1
Answer #1

We will use mathematical induction to prove the inequality.

mathbb{P}left ( igcap_{j=1}^{n} F_j ight )ge sum_{j=1}^{n}mathbb{P}(F_j) - (n-1)

For n = 1,

LHS = mathbb{P}left ( igcap_{j=1}^{1} F_j ight ) = mathbb{P}(F_1)

RHS-P(F)-(1- 1)- P(F)

Thus,  PFP) (n -1) for n = 1

and,

mathbb{P}left ( igcap_{j=1}^{n} F_j ight )ge sum_{j=1}^{n}mathbb{P}(F_j) - (n-1) holds for n = 1

Let the inequality condition hold for n = k, where k ge 1

mathbb{P}left ( igcap_{j=1}^{k} F_j ight )ge sum_{j=1}^{k}mathbb{P}(F_j) - (k-1) --- (1)

For n = k + 1

k+1 k+1

= mathbb{P}left ( igcap_{j=1}^{k} F_j ight ) + mathbb{P}(F_{k+1}) - mathbb{P}left ( igcap_{j=1}^{k} F_j igcup F_{k+1} ight )

ge sum_{j=1}^{k}mathbb{P}(F_j) - (k-1) + mathbb{P}(F_{k+1}) - mathbb{P}left ( igcap_{j=1}^{k} F_j igcup F_{k+1} ight )

= sum_{j=1}^{k}mathbb{P}(F_j) + mathbb{P}(F_{k+1}) - k+1 - mathbb{P}left ( igcap_{j=1}^{k} F_j igcup F_{k+1} ight )

= sum_{j=1}^{k+1}mathbb{P}(F_j) - (k+1-1) + left [ 1 - mathbb{P}left ( igcap_{j=1}^{k} F_j igcup F_{k+1} ight ) ight ]

As, EU k+1 is a probability value and it should lie between 0 and 1,

1 - mathbb{P}left ( igcap_{j=1}^{k} F_j igcup F_{k+1} ight ) ge 0. So,

k+1 k+1

and hence the inequality holds for n = k + 1

Hence by mathematical induction,

mathbb{P}left ( igcap_{j=1}^{n} F_j ight )ge sum_{j=1}^{n}mathbb{P}(F_j) - (n-1)